to prove :- sin3x.sin^3x+cos3x.cos^3x = cos^3*2x

to prove :-


sin3x.sin^3x+cos3x.cos^3x = cos^3*2x

Grade:12th pass

1 Answers

Aman Bansal
11 years ago

Daer Yash,
Expand Sin3x using formula :

;
sin(2x+x) = cos(2x)sin(x) + sin(2x)cos(x)
Then substitute the double angle formulae for the 2xs
= (cos^2(x) - sin^2(x))sin(x) + 2cos(x)sin(x)cos(x)
// factor out the sinx
= sin(x) (cos^2(x)-sin^2(x) + 2 cos^2(x))
=sin(x) (3cos^2(x) - sin^2(x))

and then reduce the equation to the desired form.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert